The density of mercury is 13.69 g/mL. If a thermometer contains 155 micro liters of mercury, how many atoms of mercury are in the thermometer?
1
Expert's answer
2020-09-21T06:37:13-0400

1 microliter = 1000 mL

The density of mercury "ρ = 13.69\times10^{-3}\; g/μL"

"m = V\timesρ"

"m = 155\; μL\times13.69\times10^{-3}\; g/μL = 2.12 \;g"

M(Hg) = 200.6 g/mol

"n(Hg) = \frac{m}{M} = \frac{2.12}{200.6} = 0.01 \;mol"

A mole of mercury atoms (with the formula Hg) contains 6.02×1023 Hg atoms.

Proportion:

1 mol – 6.02×1023 Hg atoms

0.01 mol – x

x = 6.02×1021 Hg atoms

Answer: 6.02×1021 Hg atoms